A car is moving along a straight line, having initial speed of v0 (v sub 0) at t = 0. The car experiences a deceleration of a = -k/v where k
is a constant and v is the speed of the car at an instant. I need to find the speed of the car as a function of time.
1
Expert's answer
2020-11-25T07:15:19-0500

We know that "a=\frac{dv}{dt}" , so we have a differential equation to solve :

"\frac{dv}{dt} = -\frac{k}{v}"

"vdv = -kdt"

We integrate both sides from "t=0" to "t=t" :

"\int_{t=0}^{t=t} vdv = -\int_0^t kdt"

"(\frac{1}{2}v^2)|^{t=t}_{t=0} = -(kt)|^{t}_0"

"\frac{1}{2}(v(t)^2-v_0^2)=-kt"

"v(t)^2=v_0^2-2kt"

So we find our final solution :

"v(t)=\pm \sqrt{v_0^2-2kt}, 0\leq t\leq \frac{v_0^2}{2k}"

We take "+" or "-" depending on the sign of "v_0" (we take the same sign that the sign of "v_0" ) , as "v(t)" must be continuous. We can also note that this solution exists only for some values of t, as for "\tau=\frac{v_0^2}{2k}, v(\tau)=0" and thus the deceleration is not well defined at "t=\tau" .